1)What is the displacement current through a circu


1)What is the displacement current through a circular loop of radiusr= 3.00 cm centered on the axis of the capacitor?


2)What is the magnitude of the magnetic field between the capacitor plates at a radiusr= 3.00 cm from the axis of the capacitor?
Answer #1

a)

Displacement current:

ID = ε0 ∂φ/∂t

φ = ∫E.dA = E A

E = Q/ε0A >>>>> φ = (Q/ε0A) A = Q/ε0

ID = ε0 ∂φ/∂t = dQ/dt = 4 C/s = 4 A

i = ID (πr2/πR2) = ID (r2/R2) = 4 * (3*3/(26*26)) = 0.0532544 ≈ 0.0533 A

b)

∫B.dA = μ0 ID

>>>> B (2πr) = μ0 i

>>>> B = (μ0 i)/2πr

>>>> B = (4*3.1415e-7)*0.0532544/(2*3.14*0.03) = 3.55 * 10^-7 T